OpenStudy (mew55):

r=1/sin(theta)+cos(theta) convert the polar equations to rectangular coordinates

OpenStudy (mew55):

\[r= \frac{ 1 }{ \sin \theta+ \cos \theta }\]

OpenStudy (anonymous):

change the equation to: \(\huge r(sin\theta+cos\theta)=1 \) now distribute: \(\huge rsin\theta+rcos\theta=1 \) rsin(theta) = ??? rcos(theta) = ???